# Variability in Footpoint Mapping of Bursty Bulk Flows Using Tsyganenko Models:

### v2

@@ 5 Discussion
Utilizing a database of 2394 busty bulk flows detected by MMS and applying four different Tsyganenko models, we mapped the BBFs to the Earth’s ionosphere. On average BBF footpoints are located in MLT coordinates around midnight, with a range of ±$\pm 3~{}\mathrm{hours}$, peaking in the pre-midnight sector (Figure 2). This location aligns with regions of intense geomagnetic activity, likely associated with the substorm activity Weigel et al. (2002).
The magnetic latitude of BBF footpoints are generally concentrated around $70^{\circ}\pm 4^{\circ}$ in both hemispheres. The Tsyganenko model T96 and T01, on average, returns a footpoint closer to the pole compared to T89. This difference could be attributed to a less stretched magnetotail in T01 and T96, as suggested by the traced field lines of the mapped BBFs (not shown). Similarly, Peredo et al. (1993) reported that the T89 model is known for excessively stretching the magnetotail in the region $-10R_{E}<X_{GSM}<-22R_{E}$.
The MLT position of the footpoints exhibits a linear relationship with the $Y_{\mathrm{GSM}}$ position of the BBF, intersection MLT=0 at $Y_{\mathrm{GSM}}=0$. This means that on average field lines represented by the Tsyganenko models do not cross the midnight meridian. Additionally, all Tsyganenko models that use $B_{y}$ and $B_{z}$ as input parameters (T96, T01, and T04) show a correlation with the IMF clock angle. For instance, figure 4 shows in the northern hemisphere that the linear regression shifts toward later MLT when $\theta_{c}\approx 270^{\circ}$ and earlier for $\theta_{c}\approx 90^{\circ}$. An opposite behavior is observed in the southern hemisphere.
The MLAT position is influenced by both the distance to Earth and the input parameter representing magnetospheric dynamics. As expected, more distant BBFs map, on average, to higher latitudes. Additionally, higher values of the magnetospheric dynamics parameter correspond to lower latitudes for the footpoints. This pattern aligns with the equatorward shift of the low-latitude boundary of the auroral oval during the growth phase of a substorm Akasofu (2017). However, the parameters used to model magnetospheric dynamics differ across the various Tsyganenko models, leading to discrepancies in the latitude of the footpoints when one parameter indicates higher magnetospheric activity than another. Interestingly, T96 shows a low correlation with the Sym-h index ($r_{p}=-0.18$ and $r_{s}=-0.08$). This is likely because, during BBF events, the Sym-h index typically reflects non-storm periods (-30 nT $<$ Sym-h $<$ 0 nT).
@@ 6 Conclusion
In this paper we provided an analysis of the variability in BBFs footpoint using four different Tsyganenko models and its impact on the number of BBF-Swarm conjunctions, with the following results:
  • Bursty bulk flows footpoints in the ionosphere are 90% of cases concentrated around $70^{\circ}\pm 4^{\circ}$ (MLAT) and around 0 MLT ±$\pm 3$ hours, with a peak in the pre-midnight sector.
  • As expected, the BBF footpoint position in MLAT coordinate is closer to the pole as the BBF distance from Earth increases and closer to the equator for larger magnitudes of the Tsyganenko input parameter that quantifies the magnetospheric current systems (i.e, Kp, Dst, $E_{kv}$, $W_{1}$ for T89, T96, T01 and T04 respectively).
  • MLAT difference between models shows footpoints closer to the pole for models T96 and T01 with respect to T89. This is related to the excessively stretched magnetotail in model T89.
  • footpoints of BBFs located at similar geocentric distances are found at lower MLAT for the model with the higher relative magnitude of the input parameter (Kp, Dst, $E_{kv}$, $W_{1}$ for T89, T96, T01 and T04 respectively) when both are normalized.
  • The difference between the footpoints computed with the T89 and T96 models are within the range of MLAT ±$\pm 2^{\circ}$, based on the 90th percentile of the data.
  • BBF footpoint position in MLT coordinate is strongly correlated with the $Y_{\mathrm{GSM}}$ position of the BBF and the interplanetary magnetic field clock angle.
  • The MLT differences between the models are within the range of ±$\pm 1$ hour, based on the 90th percentile of the data. The footpoints of the T96, T01 and T04 models at similar $Y_{\mathrm{GSM}}$ position are located at earlier MLTs compared to the T89 footpoint when the $\theta_{c}\approx 270^{\circ}$. In contrast they appear at later MLT when $\theta_{c}\approx 90^{\circ}$.
@@
  • The number of BBF-Swarm conjunctions is larger with Swarm B than Swarm A/C. Moreover, Swarm B presents, on average, more conjunctions where the footpoints distance is lower than (±$\pm 0.3$ hours in MLT and ±$\pm 0.81^{\circ}$ in MLAT) and Swarm path goes through both footpoints.
Acknowledgements.
This work was supported by the ESA 4D ionosphere initiative. ESA contract No. 383 4000143412/23/I-EB and the Swedish Research Council (Grant 2021-06259). A. P. D. received support from the Swedish National Space Agency (Grant 2020-00111). OM acknowledges support also by project MAGICS, ESA PRODEX contract 4000127660.
@@ Appendix A Application of criteria to magnetic field line mapping
In some cases, the footpoints obtained from the Tsyganenko model may be unreliable. Here, we present three different scenarios of field line tracing to illustrate this. In Figure 10a, all four Tsyganenko models (ie., T89, 96, T01 and T04) present both footpoints at the Earth’s surface and the field line makes a single loop. Figure 10b presents both footpoints at the Earth’s surface but perform several loops. Finally, 10c shows that the field line computed with T89 do not present a footpoint in the southern hemisphere.

### v3

@@ 5 Discussion
Utilizing a database of 2394 busty bulk flows detected by MMS and applying six different Tsyganenko models, we mapped the BBFs to the Earth’s ionosphere. On average, BBF footpoints are located in MLT coordinates around midnight, with a range of ±$\pm 3~{}\mathrm{hours}$, peaking in the pre-midnight sector (Figure 2). This location aligns with regions of intense geomagnetic activity, likely associated with the substorm activity Weigel et al. (2002).
The magnetic latitude of BBF footpoints are generally concentrated around $70^{\circ}\pm 5^{\circ}$ in both hemispheres. The Tsyganenko model 89, on average, returns a footpoint at lower latitudes compared with earlier models. This difference could be attributed to a more stretched magnetotail in the T89 model, as noted by Peredo et al. (1993), who reported that T89 is known to excessively stretch the magnetotail in the region $-10,R_{E}<X_{GSM}<-22,R_{E}$.
The MLT position of the footpoints exhibits a linear relationship with the $Y_{\mathrm{GSM}}$ position of the BBF, intersection MLT=0 at $Y_{\mathrm{GSM}}=0$. This means that on average the field lines represented by the Tsyganenko models do not cross the midnight meridian. Additionally, all Tsyganenko models that use $B_{y}$ and $B_{z}$ as input parameters (T96, T01, T04, T15b, and T15n) show a correlation with the IMF clock angle. For instance, Figure 4 shows in the Northern Hemisphere that the linear regression shifts toward later MLT when $\theta_{c}\approx 270^{\circ}$ and earlier for $\theta_{c}\approx 90^{\circ}$. The opposite behavior is observed in the Southern Hemisphere.
The MLAT position is influenced by both the distance to Earth and the input parameter representing magnetospheric dynamics. As expected, more distant BBFs map, on average, to higher latitudes. Additionally, higher values of the magnetospheric dynamics parameter correspond to lower latitudes for the footpoints. This pattern aligns with the equatorward shift of the low-latitude boundary of the auroral oval during the growth phase of a substorm Akasofu (2017). However, the parameters used to model magnetospheric dynamics differ across the various Tsyganenko models, leading to discrepancies in the latitude of the footpoints when one parameter indicates higher magnetospheric activity than another. Interestingly, T96 shows a low correlation with the Sym-h index ($r_{p}=-0.18$ and $r_{s}=-0.08$). This is likely because, during BBF events, the Sym-h index typically reflects non-storm periods (-30 nT $<$ Sym-h $<$ 0 nT). Similarly, the TA15 models exhibit negligible correlation between magnetic latitude and either the B-index or the N-index.
@@ 6 Conclusion
In this paper, we provided an analysis of the variability in BBFs footpoint using six different Tsyganenko models and its impact on the number of BBF-Swarm conjunctions, with the following results:
  • Based on several widely used Tsyganenko magnetic field models, 90% of the ionospheric footpoints of bursty bulk flow events inferred from MMS observations in the magnetotail are concentrated around $70^{\circ}\pm 5^{\circ}$ MLAT and 0 MLT ±$\pm 3$ hours, with a peak in the pre-midnight sector.
  • As expected, the BBF footpoint position in MLAT coordinate is closer to the pole as the BBF distance from Earth increases and closer to the equator for larger magnitudes of the Tsyganenko input parameter that quantifies the magnetospheric current systems (i.e, Kp, Dst, $E_{kv}$, $W_{1}$ for T89, T96, T01 and T04 respectively). However, the TA15 models do not exhibit this trend.
  • The difference in MLAT between models shows footpoints at lower latitudes for the T89 model. This is related to the excessively stretched magnetotail in model T89.
  • Footpoints of BBFs located at similar geocentric distances are found at lower MLAT for the model with the higher relative magnitude of the input parameter (Kp, Dst, $E_{kv}$, $W_{1}$ for the T89, T96, T01, and T04 models respectively) when both are normalized.
  • The difference between the footpoints computed with the T89 and TA15n models is within the range of MLAT ±$\pm 4^{\circ}$, based on the 90th percentile of the data.
  • The position of the BBF footpoint in the MLT coordinate is strongly correlated with the $Y_{\mathrm{GSM}}$ position of the BBF and the interplanetary magnetic field clock angle.
  • The MLT differences between the models are within the range of ±$\pm 1$ hour, based on the 90th percentile of the data. The Northern Hemisphere footpoints of the T89 model in a similar $Y_{\mathrm{GSM}}$ position are located at later MLTs compared to the TA15n footpoint when $\theta_{c}\approx 270^{\circ}$. In contrast, they appear at earlier MLT when $\theta_{c}\approx 90^{\circ}$. An opposite trend is observed when comparing (T04 and TA15n) with TA15b. For the Southern Hemisphere, these relationships are reversed.
@@
  • The number of BBF-Swarm conjunctions is larger with Swarm B than Swarm A/C, involving approximately 10% of the total database. Moreover, Swarm B presents, on average, more conjunctions where the footpoints distance is lower than (±$\pm 1$ hours in MLT and ±$\pm 4^{\circ}$ in MLAT) and Swarm path goes through both footpoints. However, the number of conjunctions is significantly reduced when considering a tighter footprint distance on the order of the BBF footpoint itself (±$\pm 0.3$ hours in MLT and ±$\pm 0.81^{\circ}$ in MLAT).
7 Acknowledgements
VL, AD, LR, SB and OM would like to acknowledge the discussions with members of ISSI Team #24-614 “Understanding ground magnetic disturbances due to field-aligned currents driven by magnetotail activity”. We thank the editor and two anonymous reviewer for evaluating this paper. We acknowledge use of NASA/GSFC’s Space Physics Data Facility’s OMNIWeb service, and OMNI data.
8 Funding
This work was supported by the ESA 4D ionosphere initiative. ESA contract No. 383 4000143412/23/I-EB and the Swedish Research Council (Grant 2021-06259). AD. received support from the Swedish National Space Agency (Grant 2020-00111). OM acknowledges support also by project MAGICS, ESA PRODEX contract 4000127660.
9 Data availability
Solar wind parameters are available at https://omniweb.gsfc.nasa.gov. Solar wind driving parameters G and W at https://rbsp-ect.newmexicoconsortium.org/data_pub/QinDenton/, parameters Bindex and Nindex at https://geo.phys.spbu.ru/~tsyganenko/empirical-models/magnetic_field/ta15. IDL Geopack DLM is available at https://korthhaus.com/idl-software/idl-geopack-dlm/. The Hp30 index is accessible through the GFZ website at https://kp.gfz.de/en/hp30-hp60/data. The Database of MMS busty bulk flows including their ionospheric footpoints computed using Tsyganenko models (T89, T96, T01, T04, TA15B, and TA15N) is available at https://doi.org/10.5281/zenodo.13789047.
@@ Appendix A Application of criteria to magnetic field line mapping
In some cases, the footpoints obtained from the Tsyganenko model may be unreliable. Here, we present three different scenarios of field line tracing to illustrate this. In Figure 10a, four Tsyganenko models (i.e., T89, 96, T01 and T04) present both footpoints at the Earth’s surface and the field line makes a single loop. Figure 10b presents both footpoints at the Earth’s surface but perform several loops. Finally, 10c shows that the field line computed with T89 does not present a footpoint in the Southern Hemisphere.